# First Comparative Exoplanetology Within a Transiting Multi-planet System: Comparing the atmospheres of V1298 Tau b and c

### v1

@@ 3.1 The transmission spectrum of V1298 Tau c
We compared the observed transmission spectrum of V1298 Tau c to a flat line and found that models with atmospheric water absorption are slightly favored (2.5). To confirm the presence of water, we conducted a free atmospheric retrieval using the open-source code PetitRadtrans (Mollière et al., 2019). We fixed the planet mass and temperature to the median of the ATMO retrieval posteriors (Table 2) and included a grey cloud deck opacity. The posterior distribution from this fit is shown in Figure 11. While we did not find constraints on CH4, our models preferred the presence of water, with a lower limit for the water volume mixing ratio () at 3 confidence level.
Comparing the spectrum of V1298 Tau c to its sibling planet, V1298 Tau b, revealed that the transit depth uncertainties for V1298 Tau c are larger by a factor of about 1.5. This higher uncertainty is likely due to including the first orbit of the visit for V1298 Tau c, known to have larger wavelength-dependent systematics. Consequently, we did not apply the common-mode correction to the spectroscopic light curves of planet c, as we did for planet b. The common-mode correction could correct for the curvature of the time-series baseline, which then allowed us to fit a linear baseline to the spectroscopic light curves of planet b.
3.2 Relative transmission spectroscopy between planet b and c
In Figure 1 (lower panel), we present the first relative transmission spectrum between two planets within the same system. This relative spectrum is independent of the star’s radius and limb darkening. Stellar inhomogeneities are known to contaminate transmission spectra (Désert et al., 2011; Rackham et al., 2019), which has also been highlighted in recent JWST observations(May et al., 2023; Moran et al., 2023; Lim et al., 2023). A relative transmission spectrum between planets theoretically bypasses stellar contamination issues, assuming consistent stellar contamination across both transits. This enables a better relative comparison of abundances between the planets without precise information on the stellar properties, including metallicity, elemental abundance ratios, and limb-darkening. However, in the case of V1298 Tau, an active star, changes in the stellar surface between observation epochs may still introduce contamination to the relative spectrum extracted at different times.
@@ 3.2 Relative transmission spectroscopy between planet b and c
| Mass [M⊕] | 7.58 | 10 (fixed) | 23 (fixed) |
3.3 Haze formation in the atmosphere of V1298 Tau planets
@@
Hazes have been found to be ubiquitous in exoplanet atmospheres (Kempton et al., 2023; Libby-Roberts et al., 2020a; Knutson et al., 2014). The temperature of V1298 Tau b and c is expected to be between 600-1000K and the host star is young and active, producing copious amounts of XUV radiation, which are important ingredients for haze formation (Yu et al., 2021).
We simulated the transmission spectrum of V1298 Tau b and c, including hazes.We assumed tholin like composition for the hazes (Khare et al., 1984). The simulations were done following the approach in Kawashima & Ikoma (2019). The details of the haze microphysics model are provided in Kawashima & Ikoma (2018). We assumed 20M⊕ and 23M⊕ (Barat et al., 2024) for planet c and b respectively. For further details see Appendix F. The simulated transmission spectrum for V1298 Tau b and c, assuming a 1% haze formation efficiency is shown in Figure 1. The effect of hazes in the HST bandpass is to damp the water absorption feature. Given the large observable water feature in V1298 Tau b’s spectrum, efficient haze formation can be ruled out for this planet. However, atmospheric models with hazes can explain the observed transmission spectrum of planet c. If we assume a lower mass for planet c (10M⊕), which is consistent with current estimates from TTV analysis, we would require even higher haze formation rates to explain the observed spectrum of V1298 Tau c.
@@ 4.1 Relative orbital distances and stellar density
In multi-planet systems, it is possible to fix one planet’s semi-major axis based on another’s measured semi-major axis, since they are related to their orbital periods ratio. However, due to known TTVs in this system, we refrain from fixing the semi-major axis in this analysis. The semi-major axis ratio from independent fits to the white light curves of V1298 Tau b and c is 2.080.1 and agrees with the expected value from their known orbital periods. Assuming circular orbits, the semi-major axis can be used to estimate stellar density (Sandford & Kipping, 2017). From the fits to our white light curves, we estimate the stellar density from the planet b and planet c to be 0.700.01gm/cm3 and 0.670.1gm/cm3. These values align with the published stellar density for V1298 Tau (0.680.1gm/cm3, David et al. (2019b)). indicating that it is inflation compared to similar-mass main sequence stars.
@@ 4.2 Limb darkening of V1298 Tau
Modeling stellar limb darkening is crucial for deriving transmission spectra. Prior studies often employed fixed limb darkening for main-sequence stars based on theoretical stellar structure models (e.g, see Deming et al., 2013; Sing et al., 2016). Barat et al. (2024) fitted a linear limb darkening coefficient for V1298 Tau, revealing significant differences compared to theoretical models for a main-sequence star with a similar spectral type (K1). Higher limb darkening compared to theoretical expectations for mature stars has previously been noted for -Pictoris (Landman et al., 2024). The fitted limb darkening coefficients and a theoretical linear limb darkening model is shown in Figure 9. This constitutes the first direct measurement of limb darkening on a weak-lined T-Tauri star.
In Section 4.1, we highlighted that V1298 Tau’s density is lower than that of similar-mass main-sequence stars. Furthermore, it has a relatively high magnetic field strength of the order of 100-300 Gauss (Finociety et al., 2023). High magnetic fields have been proposed as a potential cause for discrepancies between observed and theoretical limb darkening models (Kostogryz et al., 2024). Therefore, a combination of structural differences and higher magnetic fields compared to main-sequence stars, could explain the differences we find with the model.
While leveraging light curves from multiple planets in the same system could potentially reduce limb darkening coefficient uncertainties, the absence of ingress or egress coverage in the light curves of planet c led us to fix the limb darkening coefficients to the values from Barat et al. (2024). A comparison of the transmission spectrum of planet c derived using fixed and free limb darkening is shown in Figure 8. The transit depth uncertainties are higher by 5-10% for each spectroscopic bin in the latter case. Residuals in both cases are approximately 1.2 times the expected photon noise, indicating that stellar variability and instrumental systematics from the first orbit are likely to be the dominant sources of noise in our case.
@@ 4.3 Comparative exoplanetology between V1298 Tau b and c
A comparison between the atmospheres of V1298 Tau b and c, measured relative to atmospheric scale height (Figure 1, upper panel), shows that these two planets could host similar atmospheres. Our retrievals show that these planets are likely to have similar masses in the sub-Neptune/super-Earth regime (Table 2). Furthermore, we also find that the atmospheric metallicities of both the planets are solar/sub-solar, which is relatively low compared to their mature counterparts which have been found to have metal-rich/hazy (¿solar) atmospheres (Désert et al., 2011; Kempton et al., 2023; Madhusudhan et al., 2023; Libby-Roberts et al., 2020b). In this scenario, both the V1298 Tau planets are similar in terms of mass and composition, but their observable atmospheres are distinct from their mature counterparts.
Atmospheric hazes can be ruled out for V1298 Tau b, given the large atmospheric absorption feature, however for planet c atmospheric hazes can explain the observed spectrum.. A comparison between the transmission spectra of planet b and c shows that planet c could have a higher haze formation. Haze formation has been linked to the UV flux of the host star; a higher UV flux results in higher monomer production rates for organic hazes (Kawashima & Ikoma, 2019). Higher atmospheric metallicity could also lead to higher haze formation. The higher haze formation in planet c compared to its sibling planet could be due to difference in metallicity, temperature and incident UV flux of the two planets.

### v2

@@ 3.1 The transmission spectrum of V1298 Tau c
We ran a free atmospheric chemistry retrieval using the open-source code PetitRadtrans (Mollière et al. 2019) on the transmission spectrum of planet c. We fixed the planet mass and temperature to 17 M⊕ and 950 K, respectively. These mass and temperature values for V1298 Tau c are taken from the best fit models to the spectrum from ATMO models, described in Section 3.2. We included a grey cloud deck opacity in the free retrieval. The posterior distribution from this fit is is shown in Figure 12. Our models find water vapor in the atmosphere of planet c, with a lower limit for the water volume mixing ratio ( at 3 confidence level). We do not find any constraints on the methane abundance.
We test the robustness of the water absorption feature indicated by the free retrieval. We compared the observed transmission spectrum of V1298 Tau c (Figure 2) to a flat line using a Chi-square test. We found that models with atmospheric water absorption are slightly favored (2.5) by comparing the difference in chi-square values between a flat line and the best fit model generated using median values of the fitted parameters from the free retrieval.
We also tested the impact of fixing the planet mass and isothermal temperature, by including Gaussian priors at 950 K and 17 M⊕ with widths of 10 K and 1 M⊕. The retrieved posterior distributions and values for the water and methane abundance agree within 1 to the ones derived from the fixed mass-temperature retrievals, shown in Figure 12. Thus, we conclude that with different sets of assumptions, our retrievals suggest the presence of water vapour in the atmosphere of planet c.
The transit depth uncertainties for planet c are larger by a factor of about 1.5 compared to the spectrum of planet b (Figure 2), adopted from Barat et al. (2024). In Section 2.2, we discuss the differences in the light curve analysis between planet c and b. Since, we cannot ignore the first orbit for the visit of planet c, unlike the case for planet b visit, we cannot apply a common-mode correction to the spectroscopic light curves of planet c. Therefore, the higher uncertainty in measured transit depth for V1298 Tau c is likely due to the modelling of each spectroscopic light curve independently using RECTE to model the charge trapping and second order polynomial to model the baseline. The common-mode correction could correct for the curvature of the time-series baseline, allowing a linear baseline model as shown in the case of planet b (Barat et al. 2024).
3.2 Results from ATMO equilibrium chemistry retrievals
In this Section, we summarize the results from the equilibrium chemistry atmospheric retrievals using ATMO for both planet b and c. The ATMO models are described in Section 2.3. Figure 2 shows the best-fit atmospheric models and the observed spectra for both planets. The retrieved parameters are enlisted in Table 2, with posterior distributions shown in Figures 13, 14 for planet c and b, respectively.
We include planet mass as a free parameter for both planets in our retrieval setup. We measure a mass (with 1 confidence interval) of 17 M⊕ for planet c. The posterior distribution shown in (Figure 13) provides a 3 mass upper limit of 50 M⊕ for planet c. The measured mass of planet c from the spectrum is consistent with previous RV mass estimates for this planet: 2010 M⊕ from Sikora et al. (2023) and 50 M⊕ at 3 level of confidence (Finociety et al. 2023). Our models find sub-solar/solar atmospheric metallicity for this planet: logZ/Z⊙=-0.16. Thus, within 1 confidence interval, the metallicity of planet c is between 0.1-10 solar. We do not find any constraint on the cloud deck opacity, the equilibrium temperature or C/O ratio for planet c.
For planet b, we measure a mass of 8 M⊕ from our free mass retrievals. The mass posterior distribution for planet b (Figure 14) shows a 3 upper limit at 20 M⊕. The free mass retrieval posterior distribution from Barat et al. (2024), which used a different atmospheric model (PetitRadtrans) compared to the present study, found a mass (and 1 confidence interval) of 9 M⊕ and 3 upper limit of 23 M⊕. The median value of mass found by Barat et al. (2024) is consistent with the present results within 1, and the updated mass upper limit is marginally lower.
The discrepancy arises because of the non-isothermal T-P profile with high interior temperature and vertical mixing to explain the non-detection of methane for this planet in Barat et al. (2024). However, our models assume isothermal atmospheres, as a result retrievals favor higher isothermal temperature (800-900 K) compared to Barat et al. (2024) (600-700 K) to push the carbon chemistry into methane poor regime (Fortney et al. 2020). Therefore we can conclude that although the mass upper limits slightly differ between Barat et al. (2024) and this work, they can be considered consistent given the different model assumptions.
Our models retrieve a sub-solar atmospheric metallicity (logZ/Z⊙=-2) for planet b. This retrieved atmospheric metallicity is consistent within 1 to metallicity reported in Barat et al. (2024). We found an upper limit to the cloud opacity for this planet: log=-0.6. We do not find any constraints on the C/O ratio for planet b.
We test the impact of keeping the planet mass as a free parameter in our retrievals. To assess this, we perform retrievals by fixing the mass of both planets. For V1298 Tau c we choose two masses for this test: 10 and 50 M⊕ which are at two ends of the mass posterior distribution (Figure 13) and are also consistent with the mass upper limit from RV studies (Finociety et al. 2023). For planet b, we choose 10 and 23 M⊕. Barat et al. (2024) has shown that the HST spectrum for V1298 Tau b could be explained with a 10 M⊕ model with a highly sub-solar metallicity and higher cloud opacity, as well as with a relatively clear solar metallicity atmosphere with a planet mass of 23 M⊕. We, therefore, choose these two scenarios for this test. The posterior distributions for these fixed mass retrievals are shown in Figures 16, 15, with retrieved parameters in Table 2. The retrieved atmospheric metallicity from these fixed mass retrievals are consistent within 1 with the free mass retrieval case for both planet b and c (Table 2). For planet c, we find that increasing the mass also leads to a small increase in the retrieved atmospheric metallicity, but for all three cases (free mass, 10 M⊕ and 50 M⊕) the retrieved metallicity lies in the range 0.1-10solar for 1 confidence interval. For planet b the metallicity is sub-solar for all mass values, however metallicity increases for higher masses. Considering the 1 confidence interval, for the three cases for this planet (free mass, 10 M⊕ and 23 M⊕), the atmospheric metallicity lies within 0.01-1 solar metallicity.
From the atmospheric retrievals on V1298 Tau b and c, we conclude that both these planets have masses in the Neptune or sub-Neptune regime. The RV mass for V1298 Tau b from Suárez Mascareño et al. (2021) is inconsistent with our measurement, but the planet c mass upper limit from RVs is consistent with our mass measurement for planet c. The mass measurements of both planet b and c from this work are consistent with follow-up RV studies (Finociety et al. 2023; Sikora et al. 2023). The mass we measure for planet b is consistent with previous analysis of its HST transmission spectrum (Barat et al. 2024). Furthermore, TTV analysis of this system (Livingston et al, in prep) also find masses of both these planets in the sub-Neptune mass regime. By testing the impact of including planet mass as a free parameter in our retrievals, we conclude that the atmosphere of V1298 Tau b appears to have metallicity that is lower than or equivalent to solar composition. For V1298 Tau c, the constraints on the atmospheric metallicity are not as precise compared to planet b; however, the posterior distribution for the metallicity peaks at solar composition with 1 confidence interval between 0.1-10solar. Therefore, it is possible that both these planets are young Neptunes or sub-Neptunes in terms of their mass. But their atmospheric metallicity is much lower (at least 1 order of magnitude) than what is typically found for known mature Neptune or sub-Neptune planets (e.g. see Bean et al. 2010; Désert et al. 2011a; Miller-Ricci Kempton et al. 2012; Morley et al. 2013; Kreidberg et al. 2014; Libby-Roberts et al. 2020a; Bean et al. 2021).
@@ 3.2 Results from ATMO equilibrium chemistry retrievals
| Mass [M⊕] | 8 | 10 (fixed) | 23 (fixed) |
3.3 Haze formation in the atmosphere of V1298 Tau b and c
@@
Hazes have been found to be ubiquitous in exoplanet atmospheres (Pont et al. 2008; Désert et al. 2011a; Miller-Ricci Kempton et al. 2012; Morley et al. 2013; Kreidberg et al. 2014; Knutson et al. 2014; Libby-Roberts et al. 2020a; Kempton et al. 2023). The temperature of V1298 Tau b and c is expected to be between 600-1000K, which has been considered a sweet spot for haze formation (Brande et al. 2024). Furthermore, the host star is young and active, producing copious amounts of UV radiation, which is an important ingredient for haze formation (Kawashima & Ikoma 2018, 2019; Gao et al. 2020; Yu et al. 2021).
We simulated the transmission spectrum of V1298 Tau b and c, including hazes. We assumed tholin like composition for the hazes (Khare et al. 1984). The haze microphysics models were simulated using the models described in Kawashima & Ikoma (2018), following the approach of Kawashima & Ikoma (2019). First, using a photochemical model the abundance of gaseous species was calculated. The XUV spectrum of V1298 Tau was adopted from Duvvuri et al. (2023). Haze production rates (or efficiency of haze formation) are uncertain; therefore, we assumed a certain fraction of precursor molecules to be dissociated to form haze monomers. The haze precursor molecules are assumed to be CH4, HCN and C2H2 in our model. We derived vertical density and radius distributions of the haze particles following Kawashima & Ikoma (2018). We assumed solar elemental abundance ratio and other parameters such as eddy diffusion coefficient, monomer radius, internal density, and refractive index of haze particles similar to Kreidberg et al. (2022).
We assume 685 K and 979 K as the isothermal temperatures for these models for planet b and c, respectively (Table 1). Since the haze models are computationally expensive, we create models for two masses of planet b and c each. For planet b, we choose two masses: 10 M⊕ (consistent within 1 to the median of the ATMO retrieval mass posterior distribution for planet b) and 23 M⊕ (3 upper limit from Barat et al. (2024)). For planet c, we choose 10 M⊕, and 20 M⊕ (consistent within 1 to the median of the ATMO retrieval mass posterior distribution for planet c). We choose 10 M⊕ for planet c because from ongoing TTV studies, it is likely that the mass of planet c is in the sub-Neptune regime (Livingston et. al in prep).
In Figure 2 we show the simulated transmission spectra assuming a 1 haze production efficiency for mass of 8 and 17 M⊕ for planet b and c respectively for the ease of comparison with the ATMO models. We re-scale the planet c and b models computed with masses of 20 and 10 M⊕ respectively to 17 and 8M⊕ models, assuming that a small change in the planet mass will not significantly affect the haze models. The atmospheric signal (which is a relative measurement) is inversely proportional to the planet mass (Heng & Kitzmann 2017), through the scale height. Assuming this principle, we re-scale the haze grid models to different mass values. While significant haze formation can be ruled out for V1298 Tau b, high haze formation rates could explain the observed spectrum of V1298 Tau c.
The haze formation efficiency is an uncertain parameter in our models. Furthermore, the mass of the planets is also uncertain. So, to test the effects of this parameter on the hazy atmospheric models, we create a grid of models at different haze formation efficiency and planet mass. In Figure 3 we show the spectra of both planets with different haze formation efficiencies: 10-4 and 10-5 for planet b and 10-5 and 10-2 for planet c. For V1298 Tau b, irrespective of the mass, haze formation efficiencies as low as 10-4 can be ruled out as these do not reproduce the large water absorption feature. At 8 M⊕, a haze formation efficiency of 10-5 can potentially explain the water absorption feature for planet b, but introduces a strong slope in the continuum which is not seen in the observed spectrum. For V1298 Tau c, the haze formation rates cannot be constrained as haze formation efficiencies as low as 10-5 can explain the spectrum. However, for the 10 M⊕ for planet c we can rule out low haze formation efficiencies (10-5) at 3.5 level of confidence. Therefore, we conclude that V1298 Tau b has a haze formation efficiency lower than 10-5, whereas for V1298 Tau c the atmosphere is likely to have higher haze formation than its sibling planet.
3.4 Relative transmission spectroscopy between planet b and c
In Figure 2 (lower panel), we present the first relative transmission spectrum between two planets within the same system. By relative transmission spectrum we mean the ratio of transit depths between two planets as a function of wavelength. In principle, the relative transmission spectrum is independent of the stellar radii and limb darkening effects, assuming that both planets share similar limb darkening, i.e they transit similar parts of the stellar disk. V1298 Tau b and c have impact parameters consistent with each other within 1 (David et al. 2019b) and relatively low mutual inclination (Johnson et al. 2022). Stellar inhomogeneities are known to contaminate transmission spectra (Désert et al. 2011a; Rackham et al. 2019), which has also been highlighted in recent JWST observations(May et al. 2023; Moran et al. 2023; Lim et al. 2023). A relative transmission spectrum between multiple planets in the same system, theoretically is independent of stellar contamination, assuming similar stellar contamination across both transits. This enables a better relative comparison of abundances between the planets without precise information on the stellar properties, including metallicity, elemental abundance ratios, and limb darkening. However, in the case of V1298 Tau, an active star, changes in the stellar surface between observation epochs may still introduce contamination to the relative spectrum extracted at different times.
@@ 4.1 Relative orbital distances and stellar density
In multi-planet systems, it is possible to fix one planet’s semi-major axis based on another planet’s measured semi-major axis, since they are related to their orbital period ratio. However, due to known TTVs in this system, we refrain from fixing the semi-major axis in this analysis. The semi-major axis ratio from independent fits to the white light curves of V1298 Tau b and c is 2.080.1 and agrees with the expected value from their known orbital periods. Assuming circular orbits, the semi-major axis can be used to estimate stellar density (Sandford & Kipping 2017). From the fits to our white light curves, we estimate the stellar density from the planet b and planet c transit light curve to be 0.700.01 gm/cm3 and 0.670.1 gm/cm3. These values align with the published stellar density for V1298 Tau (0.680.1 gm/cm3, David et al. (2019a)). The relatively low density of V1298 Tau compared to similar mass main-sequence stars indicates that it is inflated and currently undergoing an evolutionary phase, and is likely to end up as a sun-like star as it approaches the main-sequence (David et al. 2019b).
@@ 4.2 Limb darkening of V1298 Tau
Modelling stellar limb darkening is crucial for deriving transmission spectra. Prior studies often used fixed limb darkening for main-sequence stars based on theoretical stellar structure models (e.g, see Sing et al. 2016). Barat et al. (2024) fitted a linear limb darkening coefficient for V1298 Tau, revealing significantly higher limb darkening compared to theoretical models for a main-sequence star with a similar spectral type (K1). The limb darkening coefficients have been derived from fitting the transit light curves of V1298 Tau b observed using HST/WFC3 G141. These observations had both ingress and egress coverage and allowed precise constraints on the limb darkening model. However, for simplicity we choose a linear limb darkening model, rather than other limb darkening models with more free parameters, since HST does not provide continuous time series. Higher limb darkening of pre-main sequence stars compared to theoretical models for mature stars has previously been noted for -Pictoris (Landman et al. 2024). The fitted limb darkening coefficients and a theoretical linear limb darkening model is shown in Figure 10. This is the first direct measurement of limb darkening on a weak-lined T-Tauri star.
In Section 4.1, we highlighted that V1298 Tau’s density is lower than that of similar-mass main-sequence stars. Furthermore, it has a relatively high magnetic field strength of the order of 100-300 Gauss (Finociety et al. 2023). High magnetic fields have been proposed as a potential cause for discrepancies between observed and theoretical limb darkening models (Kostogryz et al. 2024). Therefore, a combination of structural differences and higher magnetic fields compared to main-sequence stars, could explain the differences we find with the main-sequence limb darkening model.
While leveraging light curves from multiple planets in the same system could potentially reduce limb darkening coefficient uncertainties, the absence of ingress or egress coverage in the light curves of planet c led us to fix the limb darkening coefficients to the values from Barat et al. (2024). A comparison of the transmission spectrum of planet c derived using fixed and free limb darkening is shown in Figure 9. For the both cases, we assume a linear limb darkening law. For the free limb darkening case, we put uniform priors between 0 and 1 on the linear limb darkening coefficient. The transit depth uncertainties are higher by 5-10% for each spectroscopic bin in the latter case. Residuals in both cases are approximately 1.2 times the expected photon noise, indicating that stellar variability and instrumental systematics from the first orbit are likely to be the dominant sources of noise in our case. However, the median subtracted transit depths are consistent within 1 for all spectroscopic bins (Fig 9). For the atmospheric retrievals we use the transmission spectrum derived using the fixed limb darkening case. In Figure 9 we also show a transmission spectrum for planet c derived by fixing the linear limb darkening coefficient to the EXOCTK model shown in Figure 10. We find that the median subtracted transmission spectra for this scenario is consistent within 1 for all spectroscopic bins with the fixed limb darkening case used for the analysis in this paper.
@@ 4.3 Comparative exoplanetology between V1298 Tau b and c
In Figure 2 we show a comparison between the atmospheres of V1298 Tau b and c. Comparing the spectra in terms of the scale height of both planets (Figure 2, upper panel), shows that these planets could host atmospheres of similar nature. To compute the scale height we used the best fit mass from the ATMO posteriors for both planets (Table 2). V1298 Tau c has a smaller scale height ( by a factor of 4) compared to V1298 Tau b. Atmospheric retrievals show that the masses of both these planets are likely to be in the Neptune/sub-Neptune regime (Section 3.2). These models also show that the atmospheric metallicities of both these planets are sub-solar to solar. Mature Neptune/sub-Neptune mass planets, are known to have metal-rich/hazy (¿solar) atmospheres (Bean et al. 2010; Désert et al. 2011a; Kempton et al. 2023; Madhusudhan et al. 2023; Libby-Roberts et al. 2020b; Bean et al. 2021). Thus, V1298 Tau b and c, which are the two young Neptune/sub-Neptune progenitors, appear to have distinct atmospheres compared to their mature counterparts.
Efficient haze formation (10) can be ruled out for V1298 Tau b (Figure 3), given the large atmospheric absorption feature. However for planet c atmospheric hazes, can explain the observed spectrum. In the scenario that the mass of planet c is around 17 M⊕, we cannot constrain the haze formation efficiency for planet c, but for lower masses (10 M⊕), very low haze formation efficiency (10-5) can be ruled out. Therefore, planet c could have a higher haze formation efficiency compared to its sibling planet. Haze formation has been linked to the UV flux of the host star; a higher UV flux results in higher monomer production rates for organic hazes (Kawashima & Ikoma 2019). Higher atmospheric metallicity could also lead to higher haze formation. The higher haze formation in planet c compared to its sibling planet could be due to difference in metallicity, temperature and incident UV flux of the two planets.